OpenStudy (anonymous):

How would you use the trigonometric subtraction formula to verify this identity:

OpenStudy (anonymous):

\[\sin(\frac{ \pi }{ 2 }-x)=cosx\]

OpenStudy (zzr0ck3r):

\(\sin(\alpha-\beta)= \sin(\alpha)\cos(\beta)-\sin(\beta)\cos(\alpha)\) Using that we have \(\sin(\dfrac{\pi}{2}-x)=\sin(\dfrac{\pi}{2})\cos(x)-\sin(x)\cos(\dfrac{\pi}{2})=1*\cos(x)-\sin(x)*0=\cos(x)\)

OpenStudy (anonymous):

Can you explain to me

OpenStudy (zzr0ck3r):

which part?

OpenStudy (anonymous):

=1∗cos(x)−sin(x)∗0=cos(x)

OpenStudy (zzr0ck3r):

They said you can use this formula \[\sin(\alpha-\beta)= \sin(\alpha)\cos(\beta)-\sin(\beta)\cos(\alpha)\] If we let \(\alpha=\dfrac{\pi}{2}\) and let \(\beta=x\) we get the above result.

OpenStudy (zzr0ck3r):

well what is 1*cos(x)? what is 0*sin(x)? what is cos(x) -0?

OpenStudy (anonymous):

Yeah , but why did you add that part ?

OpenStudy (zzr0ck3r):

add?

OpenStudy (anonymous):

Like why is that part there

OpenStudy (zzr0ck3r):

I feel like there is an echo in here.